OpenStudy (anonymous):

Suppose a line has a slope of 2. If a point on the line is (5, 9), which point would also be on the line?

OpenStudy (anonymous):

formula for slope?

OpenStudy (anonymous):

y=mx+b

OpenStudy (freckles):

use point slope form: \[y-y_1=m(x-x_1)\] then replace m with your slope number replace your point (x1,y1) with the point you are given that is on the line now replace x with any value and solve for y (this will give you another point on the line)

OpenStudy (freckles):

or you could turn the point slope form into slope intercept form and use the other point as the y-intercept

OpenStudy (freckles):

either way

OpenStudy (anonymous):

@freckles can't we use slope = rise over run....so that we can directly get the point

OpenStudy (freckles):

yeah you can do that too